OpenStudy (anonymous):

Suppose that the width of a rectangle is 5 inches shorter than the length and that the perimeter of the rectangle is 80 inches. The formula for the perimeter of a rectangle is P=2L+2W.

OpenStudy (anonymous):

Let length = x and width = x - 5 So your perimeter would be: P = 2 * L + 2 * W 80 = 2*x + 2(x-5) Does that point you in the right direction?

OpenStudy (anonymous):

ummm lol I think so!

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

:D You are so welcome!

OpenStudy (anonymous):

Honestly ... does that help? Are you good with finding the solution?

OpenStudy (anonymous):

UMMM I tried it and suck lol. I have literally no idea how to figure out the rest. fml

OpenStudy (anonymous):

what I got ends up making them 10 apart instead of 5

OpenStudy (anonymous):

80 = 2*X + 2*(X -5) becomes 80 = 2X + (2X - 10) when you distribute the 2. Let's switch the equation to make it easier to read. 2X + (2X - 10) = 80 ----> 4X - 10 = 80 Add 10 to both sides to get the 4X by itself ... so 4X = 90 Divide both sides by 4 to find the value of X. X is equal to 22.5 We originally said that X is the length and X-5 is the width. So your length is 22.5 and the width is 17.5.

OpenStudy (anonymous):

OMG I love you lol.

OpenStudy (anonymous):

Oh .... you just made me so happy! I needed that today! LOL!